OpenStudy (destiny.duh):

Use a ratio to compare the values. Simplify the ratio. 30 in. to 2 ft

OpenStudy (mayankdevnani):

do you know in 1 ft how many inches are there ?

OpenStudy (destiny.duh):

12

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

\[\large \bf \frac{30~inches}{2~ft}=\frac{30~inches}{2 \times 12 ~inches}\]

OpenStudy (mayankdevnani):

find the ratio

OpenStudy (destiny.duh):

30:24

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

try to convert into simpler one

OpenStudy (destiny.duh):

ok 15:12

OpenStudy (destiny.duh):

5/4

OpenStudy (mayankdevnani):

great !

OpenStudy (destiny.duh):

thank u ^^

OpenStudy (mayankdevnani):

your welcome :)

OpenStudy (mayankdevnani):

WELCOME TO OPEN STUDY :)

OpenStudy (destiny.duh):

thank u lol